OpenStudy (anonymous):

Determine the domain of the function. f as a function of x is equal to the square root of two minus x.

OpenStudy (cwrw238):

is it f(x) = sqrt(2 - x) ?

OpenStudy (anonymous):

yes

OpenStudy (cwrw238):

the domain is the set of possible values of x Hint: as you have a square root here (2-x) cannot be negative)

OpenStudy (cwrw238):

- so for example x cannot be equal to 3 as 2-x will = -1 and there are no real square roots of -1

OpenStudy (anonymous):

So would it be x is less than or equal to 2?

OpenStudy (cwrw238):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (cwrw238):

yw

OpenStudy (cwrw238):

in interval form you would write this as (-infinity, 2]

OpenStudy (anonymous):

what about this one?? \[f(x)=\sqrt{x+3}\div(x+8)(x-2)\]

OpenStudy (cwrw238):

domain of the numerator can be found in similar way as with the first function. as for the denominator x=-8 and x = 2 cannot be part of the domain because they make the denominator zero.

OpenStudy (cwrw238):

sorry gotta go right now

OpenStudy (anonymous):

its okay! thank you!